# Unraveling the Geometry of Visual Relational Reasoning

### v1

@@ Abstract
Humans and other animals readily generalize abstract relations, such as recognizing constant in shape or color, whereas neural networks struggle. To investigate how neural networks generalize abstract relations, we introduce SimplifiedRPM, a novel benchmark for systematic evaluation. In parallel, we conduct human experiments to benchmark relational difficulty, enabling direct model-human comparisons. Testing four architectures—ResNet-50, Vision Transformer, Wild Relation Network, and Scattering Compositional Learner (SCL)—we find that SCL best aligns with human behavior and generalizes best. Building on a geometric theory of neural representations, we show representational geometries that predict generalization. Layer-wise analysis reveals distinct relational reasoning strategies across models and suggests a trade-off where unseen rule representations compress into training-shaped subspaces. Guided by our geometric perspective, we propose and evaluate SNRloss, a novel objective balancing representation geometry. Our findings offer geometric insights into how neural networks generalize abstract relations, paving the way for more human-like visual reasoning in AI.
@@ Introduction
Humans and other animals demonstrate a remarkable ability to identify and generalize abstract relations across diverse contexts. For example, recognizing the relation “constant” allows individuals to flexibly apply it to different attributes, such as a row of objects sharing the same color or shape. Furthermore, one can readily apply the same abstract relation to novel attributes never encountered before. This ability to generalize abstract relations underpins flexible reasoning, supports complex problem-solving, and enables intelligent behavior in novel situations.
A central challenge in cognitive neuroscience and artificial intelligence is to develop neural network models that can achieve similar abstract relation generalization. The Raven’s Progressive Matrices (RPM) task is a widely used benchmark for assessing abstract relational reasoning ability  (?). In this task, participants view a $3\times 3$ grid of images with the bottom-right panel missing. The goal is to identify the relational rules that govern the grid and select the correct missing panel based on these inferred rules. Inspired by this task, researchers have developed large-scale RPM-style datasets, such as Procedurally Generated Matrices (PGM)  (?) and the Relational and Analogical Visual rEasoNing (RAVEN) dataset  (?, ?), and have trained various neural architectures to solve them  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?). While many of these models achieve high test accuracy on the relational rules they are trained on, they struggle to generalize to unseen rules  (?). The absence of human behavioral data on the same generalization tasks measured in neural networks further complicates the interpretation of existing results.
This limitation raises a fundamental question: What mechanisms enable or constrain abstract relation generalization in neural networks? While prior studies have documented performance levels, the representational factors shaping generalization remain poorly understood. In particular, it is unclear how learned representations encode relational structure, how these representations differ across neural network architectures, and what geometric constraints govern generalization.
In this work, we introduce a geometric framework to systematically analyze the structure of relational representations in neural networks. To compare quantitatively network generalization performance against humans, we introduce a new benchmark, the Simplified Raven’s Progressive Matrices (SimplifiedRPM) task. Using this benchmark, we compare four representative neural networks spanning different reasoning strategies. ResNet-50  (?), Vision Transformer  (?), Wild Relation Network  (?), and Scattering Compositional Learner (SCL)  (?). We conduct human experiments, concluding that SCL aligns most closely with human relational reasoning.
SimplifiedRPM enables applying a geometric theory  (?) originally developed for generalization in visual object classification tasks to the reasoning domain. Specifically, we propose that the reasoning performance of neural networks can be attributed to the geometry of their emergent “rule manifolds”. We demonstrate that this geometric theory precisely predicts empirical generalization error in neural networks, establishing a direct link between the geometry of rule manifolds for unseen relations and generalization performance. Our analysis highlights two key geometric factors, signal and dimensionality, which interact to determine the generalizability of a network. Importantly, we observe a trade-off between signal and dimensionality, driven by a structured compression mechanism, where networks amplify the signal by constraining all class manifolds—including unseen ones—into a low-dimensional subspace spanned by the principal components of the training rules. Motivated by this theory, we introduce SNRloss, a new network training objective explicitly designed to optimize the representation geometry and show that this loss function leads to adequate generalization performance with balanced geometric features. Overall, this work provides new insights into the mechanisms underlying relational reasoning through the lens of representation geometry.
@@
To systematically assess neural networks’ ability to generalize abstract relations, we developed the Simplified Raven’s Progressive Matrices (SimplifiedRPM) task (Fig. 1A). Inspired by the original RPM task  (?), SimplifiedRPM requires models to infer a relational rule from a sample row (Fig. 1A, top) and determine which of two choice rows (Fig. 1A, bottom) follows the same rule. Each row consists of three panels with abstract objects, such as triangles or circles, arranged according to a relational rule. For example, the sample row in Fig. 1A illustrates the “constant shape” rule, where all objects share the same shape while other attributes, such as number or position, vary. SimplifiedRPM refines the original RPM task  (?) by reducing answer choices from eight to two, mitigating selection bias exploitation  (?), and presenting a single example row instead of two to remove multi-row comparisons. These changes maintain RPM’s core challenge while enabling a more focused assessment of relational reasoning.
@@
We construct a dataset for evaluating abstract relation generalization in the SimplifiedRPM task, comprising rows governed by 40 distinct relational rules. Each rule applies an abstract relation (e.g., constant) to an object attribute (e.g., shape). The abstract relations and attributes are selected to be consistent with those used in human experiments  (?, ?) while maximizing the number of rules for comprehensive evaluation. Importantly, each row corresponds to a unique rule. We generate many rows per rule, resulting in a dataset of 400k training rows, 40k validation rows, and 40k test rows. To construct a SimplifiedRPM task trial, a sample row is selected that follows one rule. A correct choice is drawn from the same rule, while an incorrect choice is drawn from a different rule. Further details are provided in the Methods section.
@@
We assess models’ ability to generalize abstract relations using a held-out rule split, where $5$ out of $40$ relational rules are excluded from training. During training, models learn relational structures from the remaining $35$ rules. For instance, a model trained on ”constant number” and ”OR shape” may never encounter ”constant shape.” After training, their performance is evaluated on the five withheld rules using the SimplifiedRPM task. This design addresses limitations in the existing RPM benchmarks. RAVEN  (?) lacks a split where some test rules are entirely unseen during training. Although PGM’s Held-out Triplet split withholds certain rules, test rows still combine seen and unseen rules, allowing models to rely on familiar rules. In contrast, SimplifiedRPM enforces a strict separation between training and test rules, ensuring a more rigorous evaluation of relational generalization. By systematically varying which rules are withheld, our benchmark provides a comprehensive assessment of models’ ability to generalize across different types of abstract relations.
@@
We evaluate four representative models on the SimplifiedRPM task: ResNet-50  (?) and ViT  (?), two benchmark vision models, and WReN  (?) and SCL  (?), which are explicitly designed for relational reasoning. These models employ different computational strategies: ResNet-50 applies a series of convolutional layers across all panels simultaneously, WReN computes pairwise panel relationships and aggregates them to infer underlying rules, ViT leverages self-attention to capture relational dependencies, and SCL employs a shared relational module to improve abstract relation recognition across object attributes. Together, these models represent diverse approaches to solving RPM tasks. All models are randomly initialized and trained end-to-end. Further details on the model architecture and hyperparameters are provided in the Methods section.
@@
The models process a row to generate a high-dimensional relational representation of its underlying relational rule (Fig. 1B). To train the models on $35$ relational rules, we appended a linear classification layer to the relational representations and optimized them end-to-end using cross-entropy loss with one-hot rule labels. Fig. 1C shows test errors across $15$ random training and held-out rules splits. All models successfully learned the relational rules, achieving test errors well below the chance level of $34/35=0.97$. Among them, SCL achieved the lowest test error, followed by ViT, WReN, and ResNet-50.
@@
We assess the trained models’ ability to generalize abstract relations to unseen rules using the SimplifiedRPM task. For each pair in the five held-out rules, we generate $500$ trials by randomly selecting a sample row that follows one rule, pairing it with a correct choice from the same rule and an incorrect choice from a different rule. In each trial, the model is sequentially presented with the sample row, followed by the two choice rows. The model computes the Euclidean distance between the relational representation of the sample row and each choice row. A trial is deemed correct if the correct choice row is closer to the sample row than the incorrect one (Fig. 1D). All models outperformed the chance generalization error of $0.5$ (Fig. 1E). Among them, SCL achieved the lowest generalization error, followed by ViT, WReN, and ResNet. Notably, SCL also outperformed the other models on the PGM benchmark’s held-out rule split  (?), aligning with our observed generalization order.
@@
Interestingly, while arithmetic rules posed the greatest challenge for models to generalize, this difficulty did not arise from poor learning of the rules. Instead, it highlights the model’s inability to generalize the arithmetic relation to new attributes. When evaluating test errors—the classification error on novel rows following the trained rules—arithmetic rules consistently had much lower errors than logical rules, which exhibited the highest errors across models (figure S1).
@@
Human performance on the original RPM task  (?) follows a hierarchy of difficulty, with progression rules the easiest and logical rules the hardest  (?, ?, ?, ?). Unlike models, humans do not receive explicit training on these rules before taking the task  (?, ?). To evaluate whether models exhibit human-like reasoning, we compare their ability to generalize to novel rules against human performance.-Among the models examined, only the SCL model replicates the human difficulty pattern across relation types (Fig. 2D). To more rigorously test this alignment, we conducted additional human studies using the SimplifiedRPM task, which allows a direct comparison of abstract relation generalization between humans and models.
@@
To encompass a range of task difficulties, we sampled $130$ trials of the SimplifiedRPM task, spanning $13$ relational rule pairs with $10$ questions per pair. We recruited $n=25$ MTurk participants, each of whom answered $60$ randomly selected questions from this pool to minimize sequential learning effects. Participants had unlimited time to respond and were compensated based on accuracy. For each trial, we recorded both accuracy and response time (see Methods for details). Human participants exhibited a mean task error of $0.270\pm 0.038$ (mean±$\pm$±s.e.m.), significantly lower than the chance level of $0.5$ (one-sample t-test, $t(12)=-5.70$, $p=4.9\times 10^{-5}$). The detailed $13$ rule pairs included in the human experiment and their corresponding human performance are in table S2.
@@
We evaluated the generalization errors of the SCL model on the same SimplifiedRPM trials used in human experiments. Specifically, we trained ten SCL models for each of the thirteen rule pairs, each using a different split of the forty rules in the dataset while ensuring that the target rule pair was always among the five held-out rules. We then measured the generalization error on the held-out pair using the same trials as those used in the human evaluation. The model achieved an average task error of $0.268\pm 0.043$ (mean±$\pm$±s.e.m.), close to the human performance. Across the thirteen rule pairs, SCL models exhibited error patterns that correlated with those observed in human participants (Spearman’s rank correlation, $\rho=0.62$, $p=0.02$) (Fig. 2E).
@@
Besides task errors, human response times—the total duration participants spent on each trial—were also strongly correlated with SCL generalization error across the thirteen rule pairs (Spearman’s rank correlation, $\rho=0.90$, $p=1.8\times 10^{-5}$) (Fig. 2F). This finding suggests that participants allocated more time to solving trials that the model found more challenging. In summary, although trained solely to classify relational rules without explicit human-like guidance, the SCL model successfully captures the human order of difficulty. In contrast, other models deviate from human performance patterns (figure S2). This alignment of SCL with human performance motivates further exploration of the mechanisms underlying how neural networks perform relational reasoning.
@@
To investigate how neural networks achieve relational reasoning and generalize their learned representations to unseen relational rules, we leveraged a geometrical theory of few-shot learning  (?). This theory provides a framework for linking errors in the Simplified RPM task to interpretable geometric properties of the relational representations.
@@
The $\text{SNR}_{a}$ depends on four key interpretable geometric terms. The first is the signal: ${\parallel\Delta\mathbf{x}_{0}\parallel^{2}=\parallel\mathbf{x}_{0}^{a}-\mathbf{x}_{0}^{b}\parallel^{2}/R_{a}^{2}}$, which measures the squared Euclidean distance between the centroids $(\mathbf{x}_{0}^{a},\mathbf{x}_{0}^{b})$ of the two rule manifolds (depicted as purple line in Fig. 3A), normalized by the mean squared radius $R_{a}^{2}$ of rule $a$. A larger signal indicates greater separation between the rule $a$ and rule $b$ manifold centroids, leading to higher $\text{SNR}_{a}$ and lower generalization error. We denote $\Delta\mathbf{x}_{0}$ as the signal direction. The second term is the bias, $R_{b}^{2}R_{a}^{-2}-1$, which captures the relative difference in the sizes of two rule manifolds. When rule manifold $a$ is larger than manifold $b$, the bias term is negative, predicting a lower $\text{SNR}_{a}$ for rule a. The third term, $D_{a}^{-1}$, represents the inverse of the dimensionality $D_{a}$ of the rule $a$ manifold. $D_{a}$, known as the ”Participation ratio”  (?), is defined as $D_{a}\equiv\left(R_{a}^{2}\right)^{2}/\sum_{i=1}^{N}\left(R_{i}^{a}\right)^{4}$. $D_{a}$ quantifies the number of dimensions along which the manifold varies significantly. Higher-dimensional manifolds are preferred for minimizing generalization errors. The last term, signal-noise overlap, is given by $\parallel\Delta\mathbf{x}_{0}\cdot\mathbf{U}_{a}\parallel^{2}$ and $\parallel\Delta\mathbf{x}_{0}\cdot\mathbf{U}_{b}\parallel^{2}$. Here, $\mathbf{U}_{a}\equiv\left[\mathbf{u}_{1}^{a}R_{1}^{a},...,\mathbf{u}_{N}^{a}R_{N}^{a}\right]$ and $\mathbf{U}_{b}\equiv\left[\mathbf{u}_{1}^{b}R_{1}^{b},...,\mathbf{u}_{N}^{b}R_{N}^{b}\right]$ are matrices of the manifold axes of variation. The signal-noise overlap quantifies the alignment between these axes of variation $(\mathbf{U}_{a},\mathbf{U}_{b})$ and the signal direction $\Delta\mathbf{x}_{0}$. Smaller overlaps result in lower generalization errors.
@@
The geometrical theory identifies three independent mechanisms that facilitate generalization: increasing signal, expanding dimensionality (D), and reducing signal-noise overlap. These mechanisms provide a structured framework for analyzing differences in generalization performance across models. Fig. 3F-I illustrates the geometry of held-out rule representations before (gray bars) and after (colored bars) training, highlighting key differences between models.
Across all models, training increases the SNR for held-out rules (Fig. 3F). When analyzing individual mechanisms, we observe that all models show an increase in signal, with SCL exhibiting a particularly pronounced rise (Fig. 3G). Model-specific differences emerge in dimensionality: ResNet, WReN, and ViT increase D, whereas SCL exhibits a decrease (Fig. 3H). For signal-noise overlap, we normalize the overlap by signal magnitude: $\frac{\parallel\Delta\mathbf{x}_{0}\cdot\mathbf{U}_{a}\parallel^{2}+\parallel\Delta\mathbf{x}_{0}\cdot\mathbf{U}_{b}\parallel^{2}}{\parallel\Delta\mathbf{x}_{0}\parallel^{2}}$. This normalization allows us to distinguish whether a high signal-noise overlap is due to a large signal magnitude or a genuine alignment between the manifold and the signal direction. We observed that ResNet, WReN, and ViT reduce overlap, whereas SCL shows an increase (Fig. 3I). In summary, ResNet, WReN, and ViT leverage all three mechanisms—enhancing signal strength, expanding D, and reducing overlap—to improve generalization. In contrast, SCL primarily relies on increasing signal strength but does so at the cost of reduced dimensionality and increased overlap.
Previously, we observed that different abstract relations—progression, arithmetic, and logical—exhibited varying levels of generalization error across models (Fig. 2A-D). To understand these differences, we examined the role of signal, dimensionality, and overlap in shaping generalization performance.
Across all models, signal strength emerged as the primary driver of generalization performance, with higher signal correlating with lower generalization error. In the SCL model, we observed a human-like generalization pattern where progression rules had significantly lower generalization errors than arithmetic and logical rules (Fig. 2D). This was explained by progression rules exhibiting the highest signal (Fig. 4D). A similar pattern was observed in other models, each exhibiting distinct generalization error trends. In ResNet, arithmetic rules’ errors were higher than logical rules due to lower signal (Fig. 4A). In WReN, arithmetic rules had the highest error, driven by the lowest signal (Fig. 4B). In ViT, progression rules had the lowest error, reflecting their highest signal (Fig. 4C).
Beyond signal, dimensionality (D) did not consistently correlate with generalization performance (Fig. 4E-H). In ResNet and ViT, progression rules had the lowest D despite their low errors, suggesting that D alone does not determine generalization ability. Signal-noise overlap also showed no significant difference across abstract relations (figure S3).
@@
Fig. 5A-D shows each model’s generalization error across layers. At the input stage, rows exhibit a generalization error of $0.475\pm 0.002$, close to the $0.5$ chance level. This suggests the rules are not yet separated at the row pixel level. Across all four models, generalization errors decrease along the layers, demonstrating progressive refinement of relational representations. Fig. 5E-H shows generalization SNR across layers, highlighting distinct relational development patterns across architectures. ResNet-50 and ViT exhibit a gradual increase in SNR, reflecting a steady accumulation of relational information (Fig. 5E,G). In contrast, WReN plateaus early at the pairwise aggregation layer, with later MLP layers failing to improve further generalization SNR (Fig. 5F). In the SCL model—which achieved the best generalization performance—we observed a sharp increase in generalization SNR at the first MLP layer (Fig. 5H).

### v2

@@ Abstract
Humans readily generalize abstract relations, such as recognizing “constant” in shape or color, whereas neural networks struggle, limiting their flexible reasoning. To investigate mechanisms underlying such generalization, we introduce SimplifiedRPM, a novel benchmark for systematically evaluating abstract relational reasoning, addressing limitations in prior datasets. In parallel, we conduct human experiments to quantify relational difficulty, enabling direct model-human comparisons. Testing four models—ResNet-50, Vision Transformer, Wild Relation Network, and Scattering Compositional Learner (SCL)—we find that SCL generalizes best and most closely aligns with human behavior. Using a geometric approach, we identify key representation properties that accurately predict generalization and uncover a fundamental trade-off between signal and dimensionality: novel relations compress into training-induced subspaces. Layer-wise analysis reveals where relational structure emerges, highlights bottlenecks, and generates concrete hypotheses about abstract reasoning in the brain. Motivated by these insights, we propose SNRloss, a novel objective explicitly balancing representation geometry. Our results establish a geometric foundation for relational reasoning, paving the way for more human-like visual reasoning in AI and opening promising avenues for extending geometric analysis to broader cognitive tasks.
@@ Introduction
Humans exhibit a remarkable ability to recognize and generalize abstract relations across diverse contexts. Once we understand an abstract relation like “constant,” we can flexibly apply it to different object attributes, such as a row of items sharing the same color, shape, or number. This generalization ability allows us to recognize patterns and make inferences in novel situations, supporting flexible reasoning and complex problem-solving.
A central challenge in both cognitive science and artificial intelligence is to understand the mechanisms that enable this kind of generalization. Raven’s Progressive Matrices (RPM) task [1] has long served as a benchmark for abstract relational reasoning. In this task, participants view a $\times 3$ 3 matrix of panels with the bottom-right panel missing. Participants must infer the relational rules governing the rows or columns, such as “constant shape across the row”, and use these inferred rules to select the missing panel from several alternatives. Decades of behavioral studies in cognitive science revealed consistent patterns of human performance on the task and have established RPM accuracy as a reliable predictor of general cognitive and problem-solving abilities  [1, 2, 3, 4, 5]. Yet, the cognitive mechanisms that support this generalization ability remain unclear.
Inspired by RPM, artificial intelligence research has developed large-scale tasks, such as PGM [6] and RAVEN [7, 8], to evaluate relational reasoning in neural networks. While models trained on these datasets [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] achieve high accuracy on learned relational rules, they generalize poorly to novel rules [19]. Recent evaluations of large-scale pre-trained vision-language models on the RAVEN task have shown that these models also struggle, despite employing advanced inference strategies such as in-context learning [20]. This limitation is further compounded by the lack of detailed human behavioral data on the same tasks, making it challenging to assess whether the model’s behavior aligns with human reasoning. While recent work has begun to explore mechanisms in simplified relational tasks, such as transitive inference [21, 22, 23] and same-different relations [24, 25, 26], broad generalization across diverse relations and realistic image inputs remains an open problem.
This gap motivates a fundamental question: What mechanisms enable or constrain the generalization of abstract relations? Specifically, how are relational rules represented in neural networks, and what geometric principles govern whether they can be flexibly applied to new contexts?
To address these questions, we introduce SimplifiedRPM, a novel benchmark designed to systematically evaluate abstract relational generalization. We assess both human performance and a range of neural network models on this benchmark. We propose that a model’s ability to generalize can be understood in terms of the geometry of its emergent rule manifolds—the representations formed for different relational rules.
To formalize this hypothesis, we extend a prior geometric framework developed for object classification [27] to the domain of relational reasoning. We show that generalization performance is accurately predicted by several measurable and interpretable rule manifold geometries. Our analysis reveals a core tradeoff: generalization is constrained by a compression mechanism that restricts both seen and unseen rule manifolds to a shared low-dimensional subspace shaped by the training rules. Motivated by these insights, we propose SNRloss, a novel training objective that directly shapes representational geometry to improve relational generalization.
Together, our findings provide a principled geometric account of abstract relation generalization, offer testable hypotheses for abstract relation reasoning in the brain, and pave the way for applying geometric analysis to broader domains in cognitive science and artificial intelligence.
Results
@@
To systematically assess neural networks’ ability to generalize abstract relations, we developed the Simplified Raven’s Progressive Matrices (SimplifiedRPM) task (Fig. 1a). Inspired by the original RPM task [1], SimplifiedRPM requires models to infer a relational rule from a sample row (Fig. 1a, top) and determine which of two choice rows (Fig. 1a, bottom) follows the same rule. Each row consists of three panels with abstract objects, such as triangles or circles, arranged according to a relational rule. For example, the sample row in Fig. 1a illustrates the “constant shape” rule, where all objects share the same shape while other attributes, such as number or position, vary. SimplifiedRPM refines the original RPM task [1] by reducing answer choices from eight to two and presenting a single example row instead of two to remove multi-row comparisons. These changes maintain RPM’s core challenge while enabling a more focused assessment of relational reasoning.
@@
We construct a dataset for evaluating abstract relation generalization in the SimplifiedRPM task, comprising rows governed by 40 distinct relational rules. Each rule applies an abstract relation (e.g., constant) to an object attribute (e.g., shape). The abstract relations and attributes are selected to be consistent with those used in human experiments [1, 4] while maximizing the number of rules for comprehensive evaluation. Importantly, each row corresponds to a unique rule. We generate many rows per rule, resulting in a dataset of 400k training rows, 40k validation rows, and 40k test rows. To construct a SimplifiedRPM task trial, a sample row is selected that follows one rule. A correct choice is drawn from the same rule, while an incorrect choice is drawn from a different rule. Further details are provided in the Methods section.
@@
We assess models’ ability to generalize abstract relations using a held-out rule split, in which 555 out of 404040 relational rules are excluded from training. During training, models learn relational structures from the remaining 353535 rules. For instance, a model trained on "constant number" and "OR shape" may never encounter "constant shape." After training, their performance is evaluated on the five withheld rules using the SimplifiedRPM task. Instead of evaluating models pertained on large-scale datasets, where the overlap between their training distribution and the tested rules is ambiguous, we train models under this controlled rule-split setup. This approach ensures precise control over the relational rules encountered during training, enabling a rigorous test of relational generalization. This addresses limitations in existing RPM benchmarks: RAVEN [7] does not include splits where test rules are entirely unseen during training, and although PGM’s Held-out Triplet split withholds some rules, test items still combine seen and unseen rules, allowing models to rely on familiar patterns. By systematically varying which rules are withheld, our benchmark provides a comprehensive assessment of models’ ability to generalize across different types of abstract relations.
@@
We evaluate four representative models on the SimplifiedRPM task: ResNet-50 [28] and ViT [29], two benchmark vision models, and WReN [6] and SCL [9], which are explicitly designed for relational reasoning. These models employ different computational strategies: ResNet-50 applies a series of convolutional layers across all panels simultaneously, WReN computes pairwise panel relationships and aggregates them to infer underlying rules, ViT leverages self-attention to capture relational dependencies, and SCL employs a shared relational module to improve abstract relation recognition across object attributes. Together, these models represent diverse approaches to solving RPM tasks. All models are randomly initialized and trained end-to-end. Further details on the model architecture and hyperparameters are provided in the Methods section.
@@
The models process a row to generate a high-dimensional relational representation of its underlying relational rule (Fig. 1b). To train the models on 353535 relational rules, we appended a linear classification layer to the relational representations and optimized them end-to-end using cross-entropy loss with one-hot rule labels. Fig. 1c shows test errors across 151515 random training and held-out rules splits. All models successfully learned the relational rules, achieving test errors well below the chance level of 34/35=0.9734/35=0.9734 / 35 = 0.97. Among them, SCL achieved the lowest test error, followed by ViT, WReN, and ResNet-50.
@@
We assess the trained models’ ability to generalize abstract relations to unseen rules using the SimplifiedRPM task. While human reasoning in RPM tasks likely involves complex cognitive processes, we follow prior work in using classification to quantify relational generalization in neural networks. This serves as a useful proxy, linking internal representations to task performance. For each pair in the five held-out rules, we generate 500500500 trials by randomly selecting a sample row that follows one rule, pairing it with a correct choice from the same rule and an incorrect choice from a different rule. In each trial, the model is sequentially presented with the sample row, followed by the two choice rows. The model computes the Euclidean distance between the relational representation of the sample row and each choice row. A trial is deemed correct if the correct choice row is closer to the sample row than the incorrect one (Fig. 1d). All models outperformed the chance generalization error of 0.50.50.5 (Fig. 1e). Among them, SCL achieved the lowest generalization error, followed by ViT, WReN, and ResNet. Notably, SCL also outperformed the other models on the PGM benchmark’s held-out rule split [19], aligning with our observed generalization order.
@@
Interestingly, while arithmetic rules posed the greatest challenge for models to generalize, this difficulty did not arise from poor learning of the rules. Instead, it highlights the model’s inability to generalize the arithmetic relation to new attributes. When evaluating test errors—the classification error on novel rows following the trained rules—arithmetic rules consistently had much lower errors than logical rules, which exhibited the highest errors across models (Supplementary Fig.1).
@@
Human performance on the original RPM task [1] follows a hierarchy of difficulty, with progression rules the easiest and logical rules the hardest [2, 3, 4, 5]. Unlike models, humans do not receive explicit training on these rules before taking the task [30, 31]. To evaluate whether models exhibit human-like reasoning, we compare their ability to generalize to novel rules against human performance. Among the models examined, only the SCL model replicates the human difficulty pattern across relation types (Fig. 2d). To more rigorously test this alignment, we conducted additional human studies using the SimplifiedRPM task, which allows a direct comparison of abstract relation generalization between humans and models.
@@
To encompass a range of task difficulties, we sampled 130130130 trials of the SimplifiedRPM task, spanning 131313 relational rule pairs with 101010 questions per pair. We focused on rules defined by attributes that humans can readily discriminate, such as object numerosity, and avoided perceptually challenging attributes, such as subtle shade differences, to reduce perceptual difficulty for human participants since the focus of this study is reasoning and not visual perception. We recruited $n=25$ MTurk participants, each of whom answered 606060 randomly selected questions from this pool to minimize sequential learning effects. Participants had unlimited time to respond and were compensated based on accuracy. For each trial, we recorded both accuracy and response time (see Methods for details). Human participants exhibited a mean task error of 0.270±0.0380.270$\pm$ 0.0380.270 ± 0.038 (mean±$\pm$±s.e.m.), significantly lower than the chance level of 0.50.50.5 (one-sample t-test, $t(12)=-5.70$, $p=4.9\times 10^{-5}$). The detailed 131313 rule pairs included in the human experiment and their corresponding human performance are in Supplementary Table 1.
@@
We evaluated the generalization errors of the SCL model on the same SimplifiedRPM trials used in human experiments. Specifically, we trained ten SCL models for each of the thirteen rule pairs, each using a different split of the forty rules in the dataset while ensuring that the target rule pair was always among the five held-out rules. We then measured the generalization error on the held-out pair using the same trials as those used in the human evaluation. The model achieved an average task error of 0.268±0.0430.268$\pm$ 0.0430.268 ± 0.043 (mean±$\pm$±s.e.m.), close to the human performance. Across the thirteen rule pairs, SCL models exhibited error patterns that correlated with those observed in human participants (Spearman’s rank correlation, $\rho=0.62$, $p=0.02$) (Fig. 2e).
@@
Besides task errors, human response times—the total duration participants spent on each trial—were also strongly correlated with SCL generalization error across the thirteen rule pairs (Spearman’s rank correlation, $\rho=0.90$, $p=1.8\times 10^{-5}$) (Fig. 2f). This finding suggests that participants allocated more time to solving trials that the model found more challenging. In summary, although trained solely to classify relational rules without explicit human-like guidance, the SCL model successfully captures the human order of difficulty. In contrast, other models deviate from human performance patterns (Supplementary Fig.2). This alignment of SCL with human performance motivates further exploration of the mechanisms underlying how neural networks perform relational reasoning.
@@
To investigate how neural networks achieve relational reasoning and generalize their learned representations to unseen relational rules, we leveraged a geometrical theory of few-shot learning [27]. This theory provides a framework for linking errors in the Simplified RPM task to interpretable geometric properties of the relational representations.
@@
The $\text{SNR}_{a}$ depends on four key interpretable geometric terms. The first is the signal: ${\parallel\Delta\mathbf{x}_{0}\parallel^{2}=\parallel\mathbf{x}_{0}^{a}-\mathbf{x}_{0}^{b}\parallel^{2}/R_{a}^{2}}$, which measures the squared Euclidean distance between the centroids $(\mathbf{x}_{0}^{a},\mathbf{x}_{0}^{b})$ of the two rule manifolds (depicted as purple line in Fig. 3a), normalized by the mean squared radius $R_{a}^{2}$ of rule $a$. A larger signal indicates greater separation between the rule $a$ and rule $b$ manifold centroids, leading to higher $\text{SNR}_{a}$ and lower generalization error. We denote $\Delta\mathbf{x}_{0}$ as the signal direction. The second term is the bias, $R_{b}^{2}R_{a}^{-2}-1$, which captures the relative difference in the sizes of two rule manifolds. When rule manifold $a$ is larger than manifold $b$, the bias term is negative, predicting a lower $\text{SNR}_{a}$ for rule a. The third term, $D_{a}^{-1}$, represents the inverse of the dimensionality $D_{a}$ of the rule $a$ manifold. $D_{a}$, known as the ”Participation ratio” [32], is defined as $D_{a}\equiv\left(R_{a}^{2}\right)^{2}/\sum_{i=1}^{N}\left(R_{i}^{a}\right)^{4}$. $D_{a}$ quantifies the number of dimensions along which the manifold varies significantly. Higher-dimensional manifolds are preferred for minimizing generalization errors. The last term, signal-noise overlap, is given by $\parallel\Delta\mathbf{x}_{0}\cdot\mathbf{U}_{a}\parallel^{2}$ and $\parallel\Delta\mathbf{x}_{0}\cdot\mathbf{U}_{b}\parallel^{2}$. Here, $\mathbf{U}_{a}\equiv\left[\mathbf{u}_{1}^{a}R_{1}^{a},...,\mathbf{u}_{N}^{a}R_{N}^{a}\right]$ and $\mathbf{U}_{b}\equiv\left[\mathbf{u}_{1}^{b}R_{1}^{b},...,\mathbf{u}_{N}^{b}R_{N}^{b}\right]$ are matrices of the manifold axes of variation. The signal-noise overlap quantifies the alignment between these axes of variation $(\mathbf{U}_{a},\mathbf{U}_{b})$ and the signal direction $\Delta\mathbf{x}_{0}$. Smaller overlaps result in lower generalization errors.
@@
The geometrical theory identifies three independent mechanisms that facilitate generalization: increasing signal, expanding dimensionality (D), and reducing signal-noise overlap. These mechanisms provide a structured framework for analyzing differences in generalization performance across models. Fig. 3f-i illustrates the geometry of held-out rule representations before (gray bars) and after (colored bars) training, highlighting key differences between models. For signal-noise overlap, we normalize the overlap by signal magnitude: $\frac{\parallel\Delta\mathbf{x}_{0}\cdot\mathbf{U}_{a}\parallel^{2}+\parallel\Delta\mathbf{x}_{0}\cdot\mathbf{U}_{b}\parallel^{2}}{\parallel\Delta\mathbf{x}_{0}\parallel^{2}}$. This normalization allows us to distinguish whether a high signal-noise overlap is due to a large signal magnitude or a genuine alignment between the manifold and the signal direction.
Across all models, training leads to increased SNR on held-out rules(Fig. 3f). Analyzing individual mechanisms reveals that SCL exhibits a pronounced increase in signal (Fig. 3g) ; however, this gain is accompanied by a decrease in dimensionality (D) and an increase in signal-noise overlap (Fig.3h,i). These opposing trends partially offset the benefits of higher signal, potentially limiting generalization. In contrast, ResNet, WReN, and ViT show more balanced geometric changes: although their signal gains are more modest, they are coupled with increased dimensionality and reduced overlap (Fig. 3g-i). These findings highlight that generalization generalization arises not from single geometric factor alone, but from the interplay of all geometric components identified by our theory.
We previously observed that generalization performance varied across abstract relation types (Fig.2a–d). To investigate the geometric basis of these differences, we analyzed the signal, dimensionality, and signal-noise overlap of representations associated with each relation type. In the SCL model, which most closely mirrored human (Fig.2d), differences in generalization were primarily governed by the signal: for example, the progression rule, which was easiest to generalize, exhibited the highest signal (Fig.4d). Similarly, in ResNet, WReN, and ViT, the arithmetic rules, with the highest generalization error, were associated with the lowest signal (Fig.4a-c). Dimensionality and overlap did not show a consistent relationship with relation type (Fig.4e–h; Supplementary Fig.3). Interestingly, arithmetic rules exhibited relatively higher dimensionality despite their poor generalization, suggesting that increased dimensionality, while potentially beneficial, is not by itself sufficient for abstraction. Together, these findings identify signal as the primary geometric correlate of generalization across abstract relations.
@@
Fig. 5a-d shows each model’s generalization error across layers. At the input stage, rows exhibit a generalization error of 0.475±0.0020.475$\pm$ 0.0020.475 ± 0.002, close to the 0.50.50.5 chance level. This suggests the rules are not yet separated at the row pixel level. Across all four models, generalization errors decrease along the layers, demonstrating progressive refinement of relational representations. Fig. 5e-h shows generalization SNR across layers, highlighting distinct relational development patterns across architectures. ResNet-50 and ViT exhibit a gradual increase in SNR, reflecting a steady accumulation of relational information (Fig. 5e,g). In contrast, WReN plateaus early at the pairwise aggregation layer, with later MLP layers failing to improve further generalization SNR (Fig. 5f). In the SCL model—which achieved the best generalization performance—we observed a sharp increase in generalization SNR at the first MLP layer, but decreased slightly at the second (final) layer (Fig. 5h).